at 65 mph how far will you travel in 5 seconds
Found 2 solutions by stanbon, opeodu:
Answer by stanbon(57203)

at 65 mph how far will you travel in 5 seconds
------
distance = rate*time
distance = (65 mph)(5 seconds) = (65 mph)((5/3600)hrs)
----
= 0.0903 miles
==============================
Cheers,
Stan H.
===========

Answer by opeodu(24)

V=65mph t=5sec(1/720h) V=d/t d=v*t=65*1/720=0.0903m